OpenStudy (anonymous):

I need help with this equation in a+bi in standard form 8-i/5+i

OpenStudy (anonymous):

\[ = 8 + 4/5\cdot i \]

OpenStudy (anonymous):

its and division line

OpenStudy (anonymous):

(3-i)/2

OpenStudy (anonymous):

can you please show the steps so I can get it, and do one on my own

OpenStudy (anonymous):

Just multiply Numetrator & Denominator with the conjugate of Denominator.

OpenStudy (anonymous):

Coniugate of den is 5-i 8-i/5+i =(8-i)(5-i)/(5+i)(5-i) =(3-i)/2

OpenStudy (anonymous):

okay thanks so much!